# Extracting the expression for the field equations of a diffeomorphism invariant theory of gravity from surface term

### v1

@@ 1 Introduction
The present paper is structured as follows. Within Section 2, the equations of motion and the surface term for diffeomorphism invariant theories of pure gravity involving the covariant derivatives of the Riemann tensor are figured out via the variation of the Lagrangians for these theories. Within Section 3, under the replacement for the variation operator by the Lie derivative along an arbitrary smooth vector field, the surface term is put into the form consisting of a symmetric rank-two tensor and an anti-symmetric one. By using an equality associated to the partial derivative of the Lagrangian density with respect to the metric, which is derived out of the Lie derivative of the Lagrangian density, then it is demonstrated that the symmetric rank-two tensor gives rise to the field equations and the anti-symmetric rank-two tensor represents the Noether charge two-form corresponding to the Noether current. Within Section 4, we give a detailed comparison between two different ways for identifying the expression for the Euler-Lagrange equations of motion with the one arising from the surface term. By the way, we obtain some equivalence relations relative to the conserved current and the expression for the field equations. Within Section 5, we reveal several relationships among three significant equations arising from the replacement of the variation operator with the Lie derivative along an arbitrary vector field in the surface term and the variation equations of the Lagrangian. Along the way, we explore the sufficient and necessary condition for the derivation of the field equations from the surface term. Within Section 6, through straightforward calculations, it is proved that the expression for the field equations originating from the surface term is divergence-free. Within Section 7, in terms of all the aforementioned analysis, the procedure for the derivation of the field equations from the surface term perspective is put forward. According to the procedure, we derive the explicit expressions for the surface term and the field equations associated to the generic theories of gravity endowed with the Lagrangian relying on the metric, the Riemann tensor and the covariant derivative operator. The last section contains our conclusions and discussions.
@@ 4 The comparison between two ways of identifying the expressions for the field equations
Apart from the above process for the handling of $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$, it has been illustrated in our previous work [20] that such an equality can be obtained from the off-shell Noether current perspective. To do so, conversely, the off-shell Noether current ${J}^{\mu}$ is adopted to produce the expression for the field equations rather than arises as its outcome. As is shown by Eq. (100), the covariant divergence of this current is associated to the identity
@@
directly originating from the variation equation (15) of the Lagrangian under the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$. As a matter of fact, the identity (52), which can be viewed as a bridge between the divergence for the conserved current and that for the expression of the field equations, constitutes the starting point to yield $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ there. By contrast, employing the identity (52) instead of Eq. (43) as the starting point is relatively simple. Besides, in that work, equation (43) is treated as the consequence resulting from $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ rather than the foundation to verify this equality like in the present work. In short, the process to obtain $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ from the off-shell Noether current perspective can be summarized as
@@
Figure 1: The equivalence relationships associated to the expression ${E}^{\mu\nu}$ for the field equations and the conserved current ${J}^{\mu}$ are unveiled. The definition (49) for the conserved current is employed. In addition, the identity (52) gets involved in establishing the equivalence between $\nabla_{\mu}{E}^{\mu\nu}=0$ and $\nabla_{\mu}{J}^{\mu}=0$.
@@ 7 The procedure for extracting the expression of the field equations out of the surface term
Step 3: In terms of the expression (25) for the surface term $\Theta^{\mu}$ with $\delta\rightarrow\mathcal{L}_{\zeta}$, under the condition that equation (2) with $\delta$ substituted by $\mathcal{L}_{\zeta}$, presented by Eq. (39), is supposed to be valid, or under the condition that equation (52) holds, the expression ${E}^{\mu\nu}$ for the field equations derived out of the variation of the Lagrangian can be built from the ${X}^{\mu\nu}$ tensor, written as ${E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}{g}^{\mu\nu}$. Here the validity for either of Eqs. (39) and (52) supplies a guarantee of the symmetry for the rank-two tensor ${X}^{\mu\nu}$. Meanwhile, the combination of ${E}^{\mu\nu}$ with the quantity $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ leads to the Noether current ${J}^{\mu}$, taking the form ${J}^{\mu}=2\zeta_{\nu}{E}^{\mu\nu}+\zeta^{\mu}{L}-\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ or ${J}^{\mu}=\nabla_{\nu}{Q}^{\mu\nu}$.
@@
Within Eq. (83), the nonnegative integer $m$ denotes the maximum number of the covariant derivative operator acting on the Riemann tensor, while the symmetric rank-two tensor $\check{S}_{\mu\nu}$ and the rank-$(i+4)$ tensors $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s are actually allowed to be arbitrary in form although they are constrained by the variation of the Lagrangian density. For convenience, as before, we impose the restriction that the last four indices in the $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$ tensor identically satisfy $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}=\Delta^{\mu\nu\rho\sigma}_{\alpha\beta\gamma\lambda}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\gamma\lambda}$ attributed to the relation $\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}=\Delta^{\alpha\beta\gamma\lambda}_{\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\alpha\beta\gamma\lambda}$. Thanks to the fact that the surface term $\Theta^{\mu}_{gen}$ is adequately determined by the second term at the right hand side of Eq. (83), it is only necessary to deal with this term regardless of the scalar $\check{S}_{\mu\nu}\delta{g}^{\mu\nu}$ for the sake of deriving the expression for the equations of motion out of the surface term $\Theta^{\mu}_{gen}$. As a consequence of doing so, the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ is connected with the divergence of the surface term $\Theta^{\mu}_{gen}$ in the way
@@
Moreover, by analogy with the derivation for the identity (LABEL:SumZiLieR), replacing the variation operator $\delta$ in the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ with the Lie derivative $\mathcal{L}_{\zeta}$ along the arbitrary smooth vector field $\zeta^{\mu}$ results in the decomposition that this scalar is written as the linear combination of a term proportional to $\mathcal{L}_{\zeta}{g}^{\mu\nu}$ with a divergence term, namely,
@@
Here $\check{W}^{\mu\nu}=\sum^{m}_{i=0}{W}^{\mu\nu}_{(i)}\big|_{{Z}\rightarrow\check{Z}}$ with the $W^{\mu\nu}_{(i)}$ tensor given by Eq. (23). In the same way, the identity (87) holds for any $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$ tensor. What is more, we observe that the pair $\big(\check{N}^{\mu\nu},\check{X}^{\mu\nu}\big)$ appearing at the right hand side of the identity (87) is non-unique. Here both the rank-two tensors $\check{N}^{\mu\nu}$ and $\check{X}^{\mu\nu}$ denote the symmetric tensor proportional to $\mathcal{L}_{\zeta}{g}^{\mu\nu}$ and the one contracted with $\zeta_{\nu}$ in the divergence term, respectively. For the sake of seeing this clearly, in addition to the pair $\big(\check{N}^{\mu\nu},\check{X}^{\mu\nu}\big)$, it is assumed that there exists another arbitrary pair for rank-two tensors $\big(\check{\mathcal{N}}^{\mu\nu},\check{\mathcal{X}}^{\mu\nu}\big)$ being independent of the vector field $\zeta^{\mu}$, in which the $\check{\mathcal{N}}^{\mu\nu}$ tensor is required additionally to be symmetric. It can be proved that a sufficient and necessary condition for
@@
As a matter of fact, it is easy to see that there are infinite number of rank-two tensors satisfying the condition (90), which accordingly bring forth an infinite number of pairs $\big(\check{\mathcal{N}}^{\mu\nu},\check{\mathcal{X}}^{\mu\nu}\big)$ fulfilling Eq. (87). However, once either of $\check{\mathcal{N}}^{\mu\nu}$ and $\check{\mathcal{X}}^{\mu\nu}$ is fixed, the other one must be determined uniquely. This implies that $\check{\mathcal{N}}^{\mu\nu}$ and $\check{\mathcal{X}}^{\mu\nu}$ are in one-to-one correspondence.
@@
with $\check{Q}^{\mu\nu}={Q}^{\mu\nu}|_{{Z}\rightarrow\check{Z}}$. Equation (92) actually provides a direct way to verify that the vector $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ can be always decomposed into the form (98). Because of the allowance for the arbitrariness of the tensor $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$, the decomposition for $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ displayed by Eq. (92) possesses an evident universality, and it merely depends upon all the $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ terms regardless of whether or not such terms stem from the variation of the Lagrangian $\sqrt{-g}{L}_{gen}$. In particular, within the situation where all the $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ terms indeed originate from the variation of the Lagrangian, that is, the sum of these terms fulfills Eq. (83), in the same spirit as the analysis performed towards the derivation of Eq. (138) within Appendix B, it is able to be verified that equation (92) can also appear as a consequence out of the coexistence for both the two equations (83) and (97) with $\delta$ substituted by $\mathcal{L}_{\zeta}$.
On the basis of the surface term (86), making use of the condition that equation (83) still holds after the variation operator $\delta$ is replaced with the Lie derivative $\mathcal{L}_{\zeta}$, namely,
@@
On the other hand, the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ has the potential to produce the expression for the equations of motion. Concretely, when this scalar is embedded into the Lie derivative of the Lagrangian density presented by Eq. (93), yielding the pair $\big(\check{\mathcal{N}}^{\mu\nu},\check{\mathcal{X}}^{\mu\nu}\big)=\big(\frac{1}{2}{L}_{gen}{g}^{\mu\nu}-\check{S}^{\mu\nu},\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\big)$, by virtue of Eqs. (89) and (91), the comparison between Eqs. (87) and (93) leads to that the rank-two symmetric and divergence-free tensor $\Omega^{\mu\nu}=\check{N}^{\mu\nu}-\check{\mathcal{N}}^{\mu\nu}=\check{X}^{\mu\nu}-\check{\mathcal{X}}^{\mu\nu}$ has to take the specific value
@@
Consequently, we obtain the expression for the field equations ${E}^{\mu\nu}_{gen}=\Omega^{\mu\nu}$ together with $\nabla_{\mu}{E}^{\mu\nu}_{gen}=0$ once again. This implies that the $\Omega^{\mu\nu}$ tensor must turn into the expression for the equations of motion if the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ is determined by the Lie derivative of the Lagrangian density. Notably, here this scalar is required to be put into the form (87), which reserves the $\check{N}_{\mu\nu}$ tensor in the term $\check{N}_{\mu\nu}\delta{g}^{\mu\nu}$ of Eq. (84) and can be interpreted as the direct consequence from replacing $\delta$ in Eq. (84) with $\mathcal{L}_{\zeta}$ by the aid of Eq. (92). Within Eq. (96), the first equality represents nothing else but the expression for the Euler-Lagrange equations of motion, while the second equality denotes the expression of the field equations arising from the surface term. As a result from Eq. (96), it is proved that both of the two expressions must coincide with each other under the condition (93).

### v2

@@ 1 Introduction
The present paper is structured as follows. Within Section 2, the equations of motion and the surface term for diffeomorphism invariant theories of pure gravity involving the covariant derivatives of the Riemann tensor are figured out via the variation of the Lagrangians for these theories. Within Section 3, under the replacement for the variation operator by the Lie derivative along an arbitrary smooth vector field, the surface term is put into the form consisting of a symmetric rank-two tensor and an anti-symmetric one. By using an equality associated to the partial derivative of the Lagrangian density with respect to the metric, which is derived out of the Lie derivative of the Lagrangian density, then it is demonstrated that the symmetric rank-two tensor gives rise to the field equations and the anti-symmetric rank-two tensor represents the Noether charge two-form corresponding to the Noether current. Within Section 4, we make a detailed comparison between two different ways for identifying the expression for the Euler-Lagrange equations of motion with the one arising from the surface term. By the way, we obtain some equivalence relations relative to the conserved current and the expression for the field equations. Within Section 5, we reveal several relationships among three significant equations arising from the replacement of the variation operator with the Lie derivative along an arbitrary vector field in the surface term and the variation equations of the Lagrangian. Along the way, we explore the sufficient and necessary condition for the derivation of the field equations from the surface term. Within Section 6, through straightforward calculations, it is proved that the expression for the field equations originating from the surface term is divergence-free. Within Section 7, in terms of all the aforementioned analysis, the procedure for the derivation of the field equations from the surface term perspective is put forward. According to the procedure, we derive the explicit expressions for the surface term and the field equations associated to the generic theories of gravity endowed with the Lagrangian relying on the metric, the Riemann tensor and the covariant derivative operator. This section encompasses the main idea of the present work. The last section contains our conclusions and discussions.
@@ 4 The comparison between two ways of identifying the expressions for the field equations
Apart from the above process for the handling of $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$, it has been illustrated in our previous work [20] that such an equality can be obtained from the off-shell Noether current perspective. To do so, conversely, the off-shell Noether current ${J}^{\mu}$ is adopted to produce the expression for the field equations rather than arises as the outcome of this expression. As is shown by Eq. (103), the covariant divergence of this current is associated to the identity
@@
directly originating from the variation equation (15) of the Lagrangian under the transformation $\delta\rightarrow\mathcal{L}_{\zeta}$. As a matter of fact, the identity (52), which can be viewed as a bridge between the divergence for the conserved current and that for the expression of the field equations, constitutes the starting point to yield $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ in the work [20]. By contrast, employing the identity (52) instead of Eq. (43) as the starting point is relatively simple. Besides, in that work, equation (43) is treated as the consequence resulting from $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ rather than the foundation to verify this equality like in the present work. In short, the process to obtain $\tilde{E}^{\mu\nu}={E}^{\mu\nu}$ from the off-shell Noether current perspective can be summarized as
@@
Figure 1: The equivalence relationships associated to the expression ${E}^{\mu\nu}$ for the field equations and the conserved current ${J}^{\mu}$ are unveiled. The definition (49) for the conserved current is employed. The expression (25) for the surface term is instrumental in establishing these equivalence relationships. Additionally, equation (52), which is equivalent to equation (39), gets involved in establishing the equivalence between $\nabla_{\mu}{E}^{\mu\nu}=0$ and $\nabla_{\mu}{J}^{\mu}=0$.
@@ 7 The procedure for extracting the expression of the field equations out of the surface term
Step 3: In terms of the expression (25) for the surface term $\Theta^{\mu}$ with $\delta\rightarrow\mathcal{L}_{\zeta}$, under the condition that equation (2) with $\delta$ substituted by $\mathcal{L}_{\zeta}$, presented by Eq. (39), is supposed to be valid, or under the condition that equation (52) holds, the expression ${E}^{\mu\nu}$ for the field equations derived out of the variation of the Lagrangian can be built from the ${X}^{\mu\nu}$ tensor, written as ${E}^{\mu\nu}={X}^{\mu\nu}-\frac{1}{2}{L}{g}^{\mu\nu}$. Here the validity for either of Eqs. (39) and (52) supplies a guarantee of the symmetry for the rank-two tensor ${X}^{\mu\nu}$, namely, ${X}^{\mu\nu}={X}^{\nu\mu}$. Meanwhile, the combination of ${E}^{\mu\nu}$ with the quantity $\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ leads to the Noether current ${J}^{\mu}$, taking the form ${J}^{\mu}=2\zeta_{\nu}{E}^{\mu\nu}+\zeta^{\mu}{L}-\Theta^{\mu}(\delta\rightarrow\mathcal{L}_{\zeta})$ or ${J}^{\mu}=\nabla_{\nu}{Q}^{\mu\nu}$.
@@
Within Eq. (83), the nonnegative integer $m$ stands for the maximum number of the covariant derivative operator acting on the Riemann curvature tensor, while the symmetric rank-two tensor $\check{S}_{\mu\nu}$ and the rank-$(i+4)$ tensors $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s are actually allowed to be arbitrary in form although they are constrained by the variation of the Lagrangian density. For convenience, like before, we impose the restriction that the last four indices in the $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$ tensor identically satisfy $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}=\Delta^{\mu\nu\rho\sigma}_{\alpha\beta\gamma\lambda}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\gamma\lambda}$ attributed to the relation $\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}=\Delta^{\alpha\beta\gamma\lambda}_{\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\alpha\beta\gamma\lambda}$. Thanks to the fact that the surface term $\Theta^{\mu}_{gen}$ is adequately determined by the second term at the right hand side of Eq. (83), it is only necessary to deal with this term regardless of the scalar $\check{S}_{\mu\nu}\delta{g}^{\mu\nu}$ for the sake of deriving the expression for the equations of motion out of the surface term $\Theta^{\mu}_{gen}$. As a consequence of doing so, the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ is connected with the divergence of the surface term $\Theta^{\mu}_{gen}$ in the way
@@
Moreover, by analogy with the derivation for the identity (133), replacing the variation operator $\delta$ in the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ with the Lie derivative $\mathcal{L}_{\zeta}$ along the arbitrary smooth vector field $\zeta^{\mu}$ results in the decomposition that this scalar is written as the linear combination of a term proportional to $\mathcal{L}_{\zeta}{g}^{\mu\nu}$ with a divergence term, namely,
@@
Here $\check{W}^{\mu\nu}=\sum^{m}_{i=0}{W}^{\mu\nu}_{(i)}\big|_{{Z}\rightarrow\check{Z}}$ with the $W^{\mu\nu}_{(i)}$ tensor given by Eq. (23). After a straightforward calculation, the covariant divergence of the $\check{X}^{\mu\nu}$ tensor turns into
$\nabla_{\mu}\check{X}^{\mu\nu}=\frac{1}{2}\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\alpha\beta\rho\sigma}\nabla^{\nu}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\alpha\beta\rho\sigma}\,.$ (89)
Equation (89) is a necessary condition for the validity of the identity (87). In the same way, the identity (87) holds for any $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$ tensor.
What is more, we observe that the pair $\big(\check{N}^{\mu\nu},\check{X}^{\mu\nu}\big)$ appearing at the right hand side of the identity (87) is non-unique. Here both the rank-two tensors $\check{N}^{\mu\nu}$ and $\check{X}^{\mu\nu}$ denote the symmetric tensor proportional to $\mathcal{L}_{\zeta}{g}^{\mu\nu}$ and the one contracted with $\zeta_{\nu}$ in the divergence term, respectively. For the sake of seeing this clearly, in addition to the pair $\big(\check{N}^{\mu\nu},\check{X}^{\mu\nu}\big)$, it is assumed that there exists another arbitrary pair for rank-two tensors $\big(\check{\mathcal{N}}^{\mu\nu},\check{\mathcal{X}}^{\mu\nu}\big)$ being independent of the vector field $\zeta^{\mu}$, in which the $\check{\mathcal{N}}^{\mu\nu}$ tensor is required additionally to be symmetric. It can be proved that a sufficient and necessary condition for
@@
As a matter of fact, it is easy to see that there are infinite number of rank-two tensors satisfying the condition (91), which accordingly bring forth an infinite number of pairs $\big(\check{\mathcal{N}}^{\mu\nu},\check{\mathcal{X}}^{\mu\nu}\big)$ fulfilling Eq. (87). However, according to Eq. (92), once either of $\check{\mathcal{N}}^{\mu\nu}$ and $\check{\mathcal{X}}^{\mu\nu}$ is fixed, the other one of them must be determined uniquely. This implies that $\check{\mathcal{N}}^{\mu\nu}$ and $\check{\mathcal{X}}^{\mu\nu}$ are in one-to-one correspondence. Therefore, within Eq. (87), when the symmetric tensor $\check{N}^{\mu\nu}$ is constrained to inherit the one proportional to the variation of the metric in Eq. (84), which is given by Eq. (85), the $\check{X}^{\mu\nu}$ tensor takes the unique form (88). Apparently, owing to the fact that the $\Omega^{\mu\nu}$ tensor is symmetric and divergence-free, it has the potential to convert to the expression for the equations of motion associated to the theories of pure gravity endowed with diffeomorphism invariance if another appropriate pair is supplied in addition to the one $\big(\check{N}^{\mu\nu},\check{X}^{\mu\nu}\big)$. This will be seen below.
@@
with $\check{Q}^{\mu\nu}={Q}^{\mu\nu}|_{{Z}\rightarrow\check{Z}}$. Like the ${Q}^{\mu\nu}$ tensor, here the anti-symmetric tensor $\check{Q}^{\mu\nu}$ is determined up to the divergence of an arbitrary three-form as well. Equation (93) actually provides a direct way to verify that the vector $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ can be always decomposed into the form (101). Because of the allowance for the arbitrariness of the tensor $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$, the decomposition for $\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})$ displayed by Eq. (93) possesses an evident universality, and it merely depends upon all the $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ terms regardless of whether or not such terms stem from the variation of the Lagrangian $\sqrt{-g}{L}_{gen}$. In particular, within the situation where each of the $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\delta\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}$ terms turns into a component arising from the variation of the Lagrangian, and the sum of all of them fulfills Eq. (83), in the same spirit as the analysis performed towards the derivation of Eq. (141) within Appendix B, it is able to be verified that equation (93) can also appear as a consequence out of the coexistence for both the two equations (83) and (100) with $\delta$ substituted by $\mathcal{L}_{\zeta}$. Additionally, within the same situation, the rank-two tensor $\check{X}^{\mu\nu}$ is symmetric. Nevertheless, the symmetry of this tensor can not be surely guaranteed in other generic cases. What is more, by the aid of Eq. (93), equation (87) is able to be interpreted as the direct consequence from replacing $\delta$ in Eq. (84) with $\mathcal{L}_{\zeta}$, which gives rise to
$\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}R_{\mu\nu\rho\sigma}=\check{N}_{\mu\nu}\mathcal{L}_{\zeta}{g}^{\mu\nu}+\nabla_{\mu}\Theta^{\mu}_{gen}(\delta\rightarrow\mathcal{L}_{\zeta})\,.$ (94)
In other words, for arbitrary tensors $\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}$s, it is verified via direct calculations that equation (84) with $\delta$ transformed into $\mathcal{L}_{\zeta}$ is valid identically.
On the basis of the surface term (86), making use of the condition that equation (83) still holds after the variation operator $\delta$ is replaced with the Lie derivative $\mathcal{L}_{\zeta}$ along the arbitrary smooth vector field $\zeta^{\mu}$, namely,
@@
Apart from the aforementioned derivation for the expression (97) relative to the equations of motion, with the help of the Lie derivative of the Lagrangian density along the vector field $\zeta^{\mu}$ presented by Eq. (95), the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ is also able to produce such an expression. Concretely, when this scalar is embedded into Eq. (95), yielding another pair
$\big(\check{\mathcal{N}}^{\mu\nu},\check{\mathcal{X}}^{\mu\nu}\big)=\left(\frac{1}{2}{L}_{gen}{g}^{\mu\nu}-\check{S}^{\mu\nu},\quad\frac{1}{2}{L}_{gen}{g}^{\mu\nu}\right)\,$, (98)
by virtue of Eqs. (90) and (92), the comparison between Eqs. (87) and (95) leads to that the rank-two symmetric and divergence-free tensor $\Omega^{\mu\nu}=\check{N}^{\mu\nu}-\check{\mathcal{N}}^{\mu\nu}=\check{X}^{\mu\nu}-\check{\mathcal{X}}^{\mu\nu}$ has to take the specific value
@@
Consequently, we obtain the expression for the field equations ${E}^{\mu\nu}_{gen}=\Omega^{\mu\nu}$ together with $\nabla_{\mu}{E}^{\mu\nu}_{gen}=0$ once again. This implies that the $\Omega^{\mu\nu}$ tensor must convert to the expression for the equations of motion if the scalar $\sum^{m}_{i=0}\check{Z}_{(i)}^{\lambda_{1}\cdot\cdot\cdot\lambda_{i}\mu\nu\rho\sigma}\mathcal{L}_{\zeta}\nabla_{\lambda_{1}}\cdot\cdot\cdot\nabla_{\lambda_{i}}{R}_{\mu\nu\rho\sigma}$ is fully determined by the Lie derivative of the Lagrangian density. Notably, here this scalar is required to be put into the form (87) in advance, which reserves the $\check{N}_{\mu\nu}$ tensor in the term $\check{N}_{\mu\nu}\delta{g}^{\mu\nu}$ of Eq. (84) and takes the alternative form (94). Within Eq. (99), the first equality represents nothing else but the expression for the Euler-Lagrange equations of motion, while the second equality denotes the expression of the field equations arising from the surface term. As a result from Eq. (99), it is proved that both of the two expressions must coincide with each other under the condition (95). This further indicates that the symmetric tensor $\check{S}^{\mu\nu}$ is in agreement with the difference between both the tensors $\check{X}^{\mu\nu}$ and $\check{N}^{\mu\nu}$, namely, $\check{S}^{\mu\nu}=\check{X}^{\mu\nu}-\check{N}^{\mu\nu}$, which becomes the identity (43) when the Lagrangian (82) is consistent with the one (1).